OpenStudy (anonymous):

Which set of line segments with the given lengths form a right triangle? 9, 27, 81 10, 24, 26 13, 15, 30 15, 18, 23

OpenStudy (anonymous):

flvs?

jimthompson5910 (jim_thompson5910):

a = 9 b = 27 c = 81 a^2 + b^2 = c^2 9^2 + 27^2 = 81^2 81 + 729 = 6561 810 = 6561 Since that last equation is false, this means that these side lengths do NOT form a right triangle

jimthompson5910 (jim_thompson5910):

Do the same with the other answer choices.

hero (hero):

until you get one where right = left

hero (hero):

then you will have found the correct one

hero (hero):

Make sure you let c equal the longest side

OpenStudy (anonymous):

really dude you guys are the best teachers but we come here to get the answer man not a teacher well okay i sound stupid haha but for real give us the answer sometimes